OpenStudy (anonymous):

ind the poler form of 1-i.hint z=r exp (θ)

OpenStudy (kainui):

So if you're trying to find the polar form of (1-i) then you have to be familiar with a couple things: \[e^{i \theta}=\cos \theta +i \sin \theta\] So this will basically give you a little line that has a length of 1 in any direction based upon the angle theta. So you can take the 1+i and notice that's 45 degrees, which is just pi/4 radians and plug that in. But remember, sine and cosine of 45 is sqrt(2)/2 so you need to multiply this by a number to get you from \[r*(\frac{ \sqrt2 }{ 2 }+ i \frac{ \sqrt2 }{ 2 })=(1+i)\]